The problem i have is that i cannot figure out what number to use the addition or subtraction rule on. How do i know?
This is the equation:
4 + 3x - 6 = 3x + 2 - x
Please help if you are willing

Answer by jessicawhidden(3)

you must first combine like terms, then get all the variables alone on one side of the equation. which way you go is up to you, as long as you end up with a postive "x".
4+3x-6 = 3x+2-x combine like terms
4-6
3x-x
3x-2 = 2x+2 get all variables alone on one side
subtract 2x from both sides
add 2 to both sides
x = 4

*after solving the equation you should be able to plug the answer back in to the origional question to check if you are right!
4+(3*4)-6= (3*4)+2-4
4+12-6 = 12+2-4
10 = 10